ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: 53C15, 53C30, 53C35 Gorodski, Claudio, University of São Paulo, Brazil Contact Sub-Riemannian Symmetric Spaces Sub-Riemannian geometry is the geometry of non-integrable k-plane distributions D on n-manifolds M, 1 < k < n, where D is equipped with a positive definite metric g. We classify the simply-connected contact sub-Riemannian symmetric spaces (these belong to a class of sub-Riemannian manifolds (M, D, g) with special symmetry properties). Reference: P. Bieliavsky, and E. Falbel, and C. Gorodski, The Classification of Simply-Connected Contact Sub-Riemannian Symmetric Spaces, to appear in Pacific J. Math. (1998).